OpenStudy (anonymous):

i dont remember what the 0 means for example -(2a)^0

OpenStudy (anonymous):

\[b^0=1\] but be careful because in your example \[-(2a)^0=-1\]

OpenStudy (anonymous):

Any number raised to the zero power is 1. So the answer is -1.

OpenStudy (anonymous):

the exponent is 0

OpenStudy (anonymous):

the old "order of operations"

OpenStudy (anonymous):

PEMDAS FTW

OpenStudy (anonymous):

so for example \[7^0=1\] but \[-7^0=-1\] and \[2\times 7^0=2\times 1=2\] etc

OpenStudy (anonymous):

caution: if a = 0, then the expression would not evaluate to 1

OpenStudy (anonymous):

okay thanks. :-)

OpenStudy (anonymous):

or -1 in your case :-P

OpenStudy (anonymous):

it is true that \[0^0\] is not a number, although there seems to be some controversy about it

OpenStudy (anonymous):

yet it approaches 1

OpenStudy (anonymous):

so it could be handy to define 0^0 as equal to 1 if it makes things easier for whatever you're trying to do.

OpenStudy (anonymous):

well.. it doesn't really approach 1 from the left... :-(